OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 4 − (x + 2) = 6 − 5(x + 4) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

4 − (x + 2) = 6 − 5(x + 4) 4-x-2=6-5x-20 -x+5x=6-4-20+2 4x=-16 x=-16/4 x=-4

OpenStudy (anonymous):

i agree with the steps of fontez its perfect